# Model-Based Diffusion Sampling for Predictive Control in Offline Decision Making

### v1

@@ Abstract
Offline decision-making requires synthesizing reliable behaviors from fixed datasets without further interaction, yet existing generative approaches often yield trajectories that are dynamically infeasible. We propose Model Predictive Diffuser (MPDiffuser), a compositional model-based diffusion framework consisting of: (i) a planner that generates diverse, task-aligned trajectories; (ii) a dynamics model that enforces consistency with the underlying system dynamics; and (iii) a ranker module that selects behaviors aligned with the task objectives. MPDiffuser employs an alternating diffusion sampling scheme, where planner and dynamics updates are interleaved to progressively refine trajectories for both task alignment and feasibility during the sampling process. We also provide a theoretical rationale for this procedure, showing how it balances fidelity to data priors with dynamics consistency. Empirically, the compositional design improves sample efficiency, as it leverages even low-quality data for dynamics learning and adapts seamlessly to novel dynamics. We evaluate MPDiffuser on both unconstrained (D4RL) and constrained (DSRL) offline decision-making benchmarks, demonstrating consistent gains over existing approaches. Furthermore, we present a preliminary study extending MPDiffuser to vision-based control tasks, showing its potential to scale to high-dimensional sensory inputs. Finally, we deploy our method on a real quadrupedal robot, showcasing its practicality for real-world control.
@@ 1 Introduction
A central challenge in decision-making is to design policies that yield behaviors which are both effective and reliable. Classical methods attempt to address this challenge through optimization, but such approaches are often limited by modeling assumptions and computational complexity (rawlings2017model). In contrast, recent work has shown that data-driven generative models can achieve the same goal by directly sampling complex behaviors from available datasets (chi2023diffusion; janner2022planning; wang2024planning; pearce2023imitating; wang2022diffusion; reuss2023goalguided; hansen2023idql; chen2021decision). These generative approaches are appealing particularly because they can capture multimodal behaviors, provide diverse candidates, and operate directly from offline data without requiring additional interaction (janner2022planning; ajay2023conditional). This raises the question of how to design effective policies when interaction is not possible.
Offline decision-making considers the problem of learning policies from previously collected datasets, without the opportunity for further interaction with the environment (Figueiredo_Prudencio_2024). This setting is particularly relevant in domains where exploration is expensive or unsafe, and where only a limited number of high-quality demonstrations are available. Recent work has shown that generative models offer a natural framework for this problem by casting policy learning as behavior synthesis—flexibly representing complex trajectory distributions and enabling diverse behavior sampling from offline data (janner2022planning; ajay2023conditional; he2023diffusion; chi2023diffusion; chen2021decision). Despite these advantages, current solutions face important limitations. In particular, they struggle to effectively leverage suboptimal data, as their generative sampling may reproduce undesirable behaviors rather than filtering them out (hester2018deep; cheng2018fast). Moreover, without explicit mechanisms to handle constraints or uncertainty, these models cannot provide reliable safety guarantees in deployment (garcia2015comprehensive).
Figure 1: Framework Overview. Left: Our proposed framework, MPDiffuser, which couples a diffusion-based planner with a diffusion-based dynamics model, complemented by a ranking module. Right: A comparison highlighting key differences between MPDiffuser and prior diffusion-based trajectory generation methods.
@@
In many real-world domains such as robotics (amodei2016concrete), healthcare (yu2021reinforcement), and autonomous driving (schwarting2018planning), policies must satisfy safety constraints in addition to achieving task goals (dulac2021challenges; garcia2015comprehensive). Enforcing safety in the offline setting is especially challenging, since constraints must be respected without further interaction. Classical safe RL methods (achiam2017constrained; tessler2018reward; fujimoto2019off; kumar2020conservative) often fail due to distribution shift and inability to validate constraints, yielding overly conservative or unsafe behavior. By contrast, model predictive control (MPC), a widely used method from classical control, enforces safety by planning over short horizons and ensuring constraint satisfaction before execution (rawlings2017model; bemporad2007robust). Inspired by this, diffusion-based trajectory generation (janner2022planning; ajay2023conditional; he2023diffusion) offers a way to produce diverse rollouts that could be checked for safety. However, existing methods generate trajectories directly in data space without enforcing system dynamics, making them unreliable when outputs are not dynamically realizable.
Contributions. Motivated by these challenges, we propose Model Predictive Diffuser (MPDiffuser), a model-based compositional framework that for offline decision making consisting of: a diffusion planner that generates diverse state–action trajectories capturing task intent, a diffusion dynamics module that refines state evolution so that sampled rollouts remain faithful to the system’s transition struc5.tures, and a ranker functioning as the final arbiter, evaluating candidate trajectories against task-specific requirements such as rewards, constraints, and specified goals. Unlike standard diffusion-based methods that rely on inverse dynamics models (ajay2023conditional; xie2025latent), our approach directly models both states and actions through two complementary diffusion processes. The dynamics model in our framework is not an inverse mapping but a forward, constraint-consistent process that enforces physical feasibility during sampling. MPDiffuser employs an alternating planner–dynamics sampling scheme that progressively balances task fidelity and feasibility throughout denoising. This procedure admits a theoretical rationale as an approximation to a distribution combining both planner priors and dynamics consistency. Consequently, MPDiffuser delivers consistent improvements in feasibility, safety, and overall decision quality across both unconstrained (D4RL) and constrained (DSRL) benchmarks. Beyond offering a more principled alternative to inverse dynamics–based approaches, our compositional formulation with a forward dynamics diffusion model unlocks distinctive capabilities: it achieves higher sample efficiency by exploiting even low-quality data for dynamics learning, adapts rapidly to variations in system dynamics, and seamlessly integrates diverse objectives and constraints through candidate ranking—a common approach in diffusion-based control, yet one whose reliability is contingent on feasibility of the generated sequences. We further demonstrate that MPDiffuser scales to visual domains and outperforms existing diffusion-based approaches on a preliminary study. Finally, we demonstrated practicality of MPDiffuser through deployment on a quadrupedal robot.
@@ 2.1 Problem Setup
We consider a finite-horizon constrained Markov decision process (CMDP) defined by the tuple $(\mathcal{X},\mathcal{U},P,r,c,d,T),$ where $\mathcal{X}$ is the state space, $\mathcal{U}$ the action space, $P(x^{\prime}\mid x,u)$ the transition kernel, $r:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ the reward function, $c:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{+}^{m}$ a vector of costs, $d\in\mathbb{R}_{+}^{m}$ the available cost budget, and $T$ the horizon. The objective is to derive a policy $\pi$ that maximizes expected cumulative reward while respecting cost-budget constraints:
$\max_{\pi}\;\mathbb{E}\!\left[\sum_{t=0}^{T-1}r(x_{t},u_{t})\right]\quad\text{s.t.}\quad\mathbb{E}\!\left[\sum_{t=0}^{T-1}c_{j}(x_{t},u_{t})\right]\leq d_{j},\;\;j=1,\dots,m.$ (1)
In the online setting, the problem may be directly addressed by interacting with the environment to estimate value functions or optimize the policy. In many practical scenarios, however, such interaction is costly, unsafe, or altogether unavailable. Instead, we adopt the offline setting, where learning must proceed entirely from a fixed dataset $\mathcal{D}=\{\xi_{i}\}_{i=1}^{N}$ of existing trajectories $\xi_{i}=\{(x_{t},u_{t},r(x_{t},u_{t}),c(x_{t},u_{t}))\}_{t=0}^{T}$.
Trajectories in $\mathcal{D}$ may come from one or multiple behavior policies that are suboptimal or unsafe. The objective in offline constrained decision making is to learn a policy that maximizes reward while satisfying cost budgets, without further interaction. A common approach is to extend value-based methods such as $Q$-learning to jointly estimate reward and cost value functions, and then optimize a policy (lee2022coptidice). However, with fixed data, value estimates degrade outside the dataset’s support: as the policy departs from the behavior policies that generated $\mathcal{D}$, it induces trajectories poorly represented in the data, causing compounding errors and potential constraint violations.
An alternative viewpoint is to focus directly on synthesizing state-action trajectories. Since rewards and costs depend on how actions drive the system’s evolution, full state–action rollouts provide a natural mechanism for evaluating task objectives and constraints. This trajectory-level perspective avoids unstable extrapolation, while offering a principled way to compute cumulative returns and costs. It therefore motivates generative approaches that explicitly model state–action trajectories.
@@ 2.2 Trajectory Generation with Diffusion Models
As introduced by sohl2015deep and refined by ho2020denoising, diffusion models are a class of generative models that approximate complex data distributions by reversing a gradual noising process. Due to their success various domains, they have recently been applied to sequential decision-making, where the objects of interest are state–action trajectories $\tau=(x_{1:H},u_{0:H-1})$ of horizon $H$. By fitting a diffusion model to offline trajectories, one can approximate the conditional distribution $p_{\theta}(\tau\mid x_{0})$ and sample rollouts that closely resemble the dataset (janner2022planning).
The forward process incrementally perturbs a trajectory with Gaussian noise:
$q_{k\mid k-1}(\tau^{k}\mid\tau^{k-1})$ $=\mathcal{N}\!\left(\tau^{k};\sqrt{1-\beta_{k}}\,\tau^{k},\;\beta_{k}I\right),\;q_{k\mid 0}\coloneqq\mathcal{N}(\tau^{k};\sqrt{\bar{\alpha}_{k}},(1-\bar{\alpha}_{k})I$ (2)
where the variance schedule $\{\beta_{k}\}_{k=0}^{K}$ is fixed in advance and $\bar{\alpha}_{k}$ is defined by $\beta_{k}$. Accordingly, the reverse process seeks to undo this corruption using a score function, defined as:
$s_{\theta}(\tau^{k},k)\;\approx\;\nabla_{\tau^{k}}\log q_{k}(\tau^{k}),\quad q_{k}(\tau^{k}):=\int q_{k\mid 0}(\tau^{k}\mid\tau^{0})\,p_{\text{data}}(\tau^{0})\,d\tau^{0}.$
@@
$p_{k-1\mid k}(\tau^{k-1}\mid\tau^{k})$ $=\mathcal{N}\left(\tau^{k-1};\frac{1}{\sqrt{\alpha_{k}}}\Big(\tau^{k}+\beta_{k}\,s_{\theta}(\tau^{k},k)\Big),\sigma_{k}^{2}I\right),$ (3)
@@
$\mathcal{L}(\theta)=\mathbb{E}_{k,\;\tau,\;\epsilon}\big[\;\|\epsilon-\epsilon_{\theta}(\tau^{k},k)\|^{2}\;\big].$
This objective reduces to scaled score matching, ensuring that $\epsilon_{\theta}$ implicitly learns the score function.
@@
In many applications, reproducing typical trajectories is not sufficient: we often require rollouts that achieve high return and respect constraints. This motivates conditional trajectory generation, where the model learns $p_{\theta}(\tau\mid x_{0},y)$ for some condition $y(\tau)$, such as a target return or cost budget. A practical mechanism for enforcing such conditions is classifier-free guidance (ho2022classifier), which combines unconditional and conditional noise predictors:
$\hat{\epsilon}=\epsilon_{\theta}(\tau^{k},\emptyset,k)+\omega\Big(\epsilon_{\theta}(\tau^{k},y(\tau),k)-\epsilon_{\theta}(\tau^{k},\emptyset,k)\Big),$ (4)
where $\emptyset$ denotes a fixed null input value for the condition and $\omega>0$ controls the guidance strength. This yields trajectory samples aligned with task objectives while retaining coverage of the dataset. To train both pathways in a single model, one uses conditional dropout: the condition $y(\tau)$ is randomly masked with probability $p$, controlled by a Bernoulli variable $\beta$:
$\mathcal{L}(\theta)=\mathbb{E}_{k,\;\tau,\;\epsilon,\;\beta\sim\mathrm{Bern}(p)}\Big[\;\|\epsilon-\epsilon_{\theta}(\tau^{k},(1-\beta)\,y(\tau)+\beta\,\emptyset,k)\|^{2}\;\Big].$
Here $\beta=1$ masks the condition, training both a conditional and unconditional predictor simultaneously. At inference, the two are recombined via classifier-free guidance (see eq. equation 4).
Figure 2: Illustrative scenario: We compare sampled state trajectories with the open-loop simulations obtained by simulating the sampled actions on a 5-dimensional kinematic bike model. Diffuser fails to generate admissible state trajectories that reach the goal, Decision Diffuser produces plausible states whose actions diverge under simulation. In contrast, MPDiffuser yields trajectories that remain faithful to the system dynamics.
@@ 3 Method
In the following, we introduce the components of our compositional framework, describe the sampling procedure, and demonstrate how it can be used for the constrained decision-making problem.
@@ 3.1 Framework Components
$\mathcal{L}^{\mathrm{pl}}(\theta)=\mathbb{E}_{\tau,k,\;\epsilon_{x},\epsilon_{u}\;\beta\sim\mathrm{Bern}(p)}\left[\;\big\|[\epsilon_{x},\epsilon_{u}]-\epsilon_{\theta}^{\mathrm{pl}}\!\big(\tau^{k}_{x},\tau_{u}^{k}\mid k,x_{0},(1-\beta)\,y(\tau)+\beta\,\emptyset\big)\big\|^{2}\;\right].$ (5)
@@
Complementing the planner, the dynamics model acts as the feasibility filter—refining state trajectories so that imagined rollouts remain faithful to the system’s underlying physics. In contrast to the planner, which models full state–action rollouts, the dynamics model is a conditional diffusion model over states only, given the initial state $x_{0}$, an action sequence with corrputed by the same noise level as states $\tau_{u}^{k}$, and conditioning vector $y(\tau)$. By treating actions purely as inputs, the model dedicates its capacity to capturing state transitions, thereby yielding sharper dynamics consistency than the planner. We train our dynamics diffusion model $\epsilon_{\vartheta}^{\mathrm{dyn}}(\tau_{x}^{k}\mid\tau_{u}^{k},k,x_{0},y(\tau))$ using:
$\mathcal{L}^{\mathrm{dyn}}(\vartheta)=\mathbb{E}_{\tau,k,\;\epsilon_{x},\epsilon_{u}\;\beta\sim\mathrm{Bern}(p)}\left[\;\big\|\epsilon_{x}-\epsilon_{\vartheta}^{\mathrm{dyn}}\!\big(\tau_{x}^{k}\mid\tau_{u}^{k},k,x_{0},(1-\beta)\,y(\tau)+\beta\,\emptyset\big)\big\|^{2}\;\right].$ (6)
@@
The ranker is a practical task- and safety-aware module that evaluates sampled trajectories against desired criteria, selecting among rollouts those that achieve high reward, respect safety budgets. Within our framework, it is treated as a flexible scoring function $\rho(\tau)$ that assigns preference values to trajectories. This allows incorporation of both domain knowledge and data-driven objectives. Formally, given sampled trajectories $\{\tau_{j}\}$, the ranker outputs $\tau^{\star}=\arg\max_{\tau_{j}}\rho(\tau_{j})$, with $\rho$ defined by the task e.g. maximizing return under constraints, or minimizing goal distance. This design balances flexibility and structure: when objectives are clearly specified, $\rho$ can be explicitly defined analytically, while in settings with implicit preferences, $\rho$ may be learned from data.

### v2

@@ Abstract
Offline decision-making via diffusion models often produces trajectories that are misaligned with system dynamics, limiting their reliability for control. We propose Model Predictive Diffuser (MPDiffuser), a compositional diffusion framework that combines a diffusion planner with a dynamics diffusion model to generate task-aligned and dynamically plausible trajectories. MPDiffuser interleaves planner and dynamics updates during sampling, progressively correcting feasibility while preserving task intent. A lightweight ranking module then selects trajectories that best satisfy task objectives. The compositional design improves sample efficiency and adaptability by enabling the dynamics model to leverage diverse and previously unseen data independently of the planner. Empirically, we demonstrate consistent improvements over prior diffusion-based methods on unconstrained (D4RL) and constrained (DSRL) benchmarks, and validate practicality through deployment on a real quadrupedal robot.
Machine Learning, ICML
@@ 1 Introduction
A central challenge in decision-making is to design policies that yield behaviors which are both effective and reliable. Classical approaches address this challenge through optimization, but are often constrained by modeling assumptions and computational complexity (Rawlings et al., 2017). Conversely, recent work has shown that data-driven generative models can achieve the same goal by sampling complex behaviors from available data (Chi et al., 2023; Janner et al., 2022; Wang et al., 2024; Pearce et al., 2023; Wang et al., 2023; Reuss et al., 2023; Chen et al., 2021; Ajay et al., 2023). These methods are particularly appealing in settings where interaction is costly or unavailable (Janner et al., 2022).
Offline decision-making formalizes this setting, requiring policies to be learned solely from previously collected data without further interaction (Figueiredo Prudencio et al., 2024). While generative models provide a flexible framework for synthesizing candidate trajectories in this regime, existing approaches face important limitations. In particular, they struggle to effectively leverage suboptimal or heterogeneous data, often reproducing undesirable behaviors present in the dataset (Hester et al., 2018; Cheng et al., 2018). Moreover, without explicit mechanisms to enforce dynamics consistency or constraints, these methods offer limited reliability and weak safety guarantees at deployment time (Garcia and Fernandez, 2015).
@@
In many real-world domains, including robotics (Amodei et al., 2016), healthcare (Yu et al., 2021a), and autonomous driving (Schwarting et al., 2018), policies must satisfy safety constraints in addition to achieving task goals (Dulac-Arnold et al., 2021; Garcia and Fernandez, 2015). Enforcing safety offline is particularly challenging, as constraints must be satisfied without further interaction. Classical safe RL methods (Achiam et al., 2017; Tessler et al., 2018; Fujimoto et al., 2019; Kumar et al., 2020) often fail under distribution shift, leading to conservative or unsafe behavior. In contrast, classical control methods enforce safety through short-horizon planning with explicit constraints (Bemporad and Morari, 2007; Rawlings et al., 2017). In this spirit, diffusion-based trajectory generation (Janner et al., 2022; Ajay et al., 2023) provides a natural mechanism for producing diverse candidate rollouts, but existing methods operate directly in data space without enforcing system dynamics, limiting the reliability of the sampled trajectories.
Figure 1: Framework Overview. Left: MPDiffuser, which couples a diffusion planner with a diffusion dynamics model, complemented by a ranking module. Right: Comparison highlighting key differences between our method and prior trajectory-level diffusion methods.
Contributions. Motivated by these challenges, we propose Model Predictive Diffuser (MPDiffuser), a model-based compositional framework for offline decision making that combines three components: (i) a diffusion planner that generates diverse, task-aligned trajectories; (ii) a diffusion dynamics model that refines states to enforce consistency with system dynamics; and (iii) a ranker that selects trajectories satisfying task-specific objectives and constraints.
MPDiffuser employs an alternating sampling scheme in which task-aligned proposals are repeatedly corrected by a diffusion dynamics model during sampling. This design balances task fidelity with dynamics feasibility and admits a theoretical interpretation as approximating a distribution combining planner priors with dynamics consistency. In contrast to prior diffusion methods that rely on inverse dynamics models (Ajay et al., 2023) or post-hoc forward models (Zhou et al., 2025), MPDiffuser directly models both states and actions, with the dynamics model acting as an active, dynamics aligned component of the sampling process.
Empirically, MPDiffuser achieves consistent improvements in feasibility, safety, and decision quality across both unconstrained (D4RL) and constrained (DSRL) benchmarks. The compositional design improves sample efficiency by allowing the dynamics model to exploit low-quality and heterogeneous data, enables rapid adaptation to changes in system dynamics, and supports flexible integration of objectives and constraints via trajectory ranking. We further demonstrate scalability to visual domains and validate practical deployment on a quadrupedal robot.
@@ 2.1 Problem Setup
We consider a finite-horizon constrained Markov decision process (CMDP) defined by the tuple $(\mathcal{X},\mathcal{U},P,r,c,d,T),$ where $\mathcal{X}$ is the state space, $\mathcal{U}$ the action space, $P(x^{\prime}\mid x,u)$ the transition kernel, $r:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ the reward function, $c:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{+}^{m}$ a vector of costs, $b\in\mathbb{R}_{+}^{m}$ the available cost budget, and $T$ the horizon. The objective is to derive a policy $\pi$ that maximizes expected cumulative reward while respecting cost constraints:
$\max_{\pi}\mathbb{E}\!\left[\sum_{t=0}^{T-1}r(x_{t},u_{t})\right]\text{ s.t. }\mathbb{E}\!\left[\sum_{t=0}^{T-1}c_{j}(x_{t},u_{t})\right]$ $\leq b_{j},$
$j$ $\in\mathbb{I}_{[1,m]}.$
We consider the offline setting, where interaction with the environment is not available and learning proceeds from a fixed dataset $\mathcal{D}=\{\xi_{i}\}_{i=1}^{N}$ of trajectories, with $\xi_{i}=\{(x_{t},u_{t},r(x_{t},u_{t}),c(x_{t},u_{t}))\}_{t=0}^{T}$. The available trajectories may be collected from multiple policies and can be suboptimal or unsafe. The objective is to learn a policy that maximizes cumulative reward while satisfying cost budgets.
A common approach extends value-based methods to jointly estimate reward and cost value functions and optimize a policy (Lee et al., 2022). With fixed data, however, value estimates deteriorate outside the dataset’s support: as the learned policy deviates from the behavior policies in the dataset, it induces poorly represented trajectories, leading to compounding errors and constraint violations.
An alternative viewpoint is to focus directly on synthesizing trajectories. Since rewards and costs depend on how actions drive the system’s evolution, full state–action rollouts provide a natural mechanism for evaluating task objectives and constraints. This trajectory-level perspective avoids unstable extrapolation, while offering a principled way to compute returns and costs. It therefore motivates generative approaches that explicitly model state–action trajectories.
@@ 2.2 Trajectory Generation with Diffusion Models
As introduced by Sohl-Dickstein et al. (2015) and refined by Ho et al. (2020), diffusion models are a class of generative models that approximate complex data distributions by reversing a gradual noising process. Due to their success in various domains, they have recently been applied to decision-making, where the objects of interest are state–action trajectories $\tau=(x_{1:H},u_{0:H-1})$ of horizon $H$. By learning a diffusion model from the data, one can approximate the conditional distribution $p_{\theta}(\tau\mid x_{0})$ and sample rollouts that resemble the dataset (Janner et al., 2022).
The forward process incrementally perturbs a trajectory:
$q_{k\mid k-1}(\tau^{k}\mid\tau^{k-1})$ $=\mathcal{N}\!\left(\tau^{k};\sqrt{1-\beta_{k}}\,\tau^{k},\;\beta_{k}I\right),$
$q_{k\mid 0}$ $=\mathcal{N}(\tau^{k};\sqrt{\bar{\alpha}_{k}},(1-\bar{\alpha}_{k})I$
where the variance schedule $\{\beta_{k}\}_{k=0}^{K}$ is fixed in advance and $\bar{\alpha}_{k}$ is defined by $\beta_{k}$. Accordingly, the reverse process seeks to undo this corruption using the score function:
$s_{\theta}(\tau^{k},k)$ $\;\approx\;\nabla_{\tau^{k}}\log q_{k}(\tau^{k}),$
$q_{k}(\tau^{k})$ $=\int q_{k\mid 0}(\tau^{k}\mid\tau^{0})\,p_{\text{data}}(\tau^{0})\,d\tau^{0}.$
@@
$p_{k-1\mid k}(\tau^{k-1}\mid\tau^{k})=\mathcal{N}\left(\tau^{k-1};\frac{\tau^{k}}{\sqrt{\alpha_{k}}}+\frac{\beta_{k}}{\sqrt{\alpha_{k}}}\,s_{\theta}(\tau^{k},k),\sigma_{k}^{2}I\right),$
@@
$\mathcal{L}(\theta)=\mathbb{E}_{\tau,k,\epsilon}\big[\;\|\epsilon-\epsilon_{\theta}(\tau^{k},k)\|^{2}\;\big],$
which reduces to score matching under a simple reparameterization, ensuring that $\epsilon_{\theta}$ learns the score function.
@@
In many applications, reproducing typical trajectories is not sufficient: we often require rollouts that are task-aligned. This motivates conditional trajectory generation, where the model learns $p_{\theta}(\tau\mid x_{0},y)$ for some condition $y(\tau)$, such as a target return or cost budget. A practical mechanism for enforcing such conditions is classifier-free guidance (Ho and Salimans, 2021), which combines unconditional and conditional noise predictors:
$\hat{\epsilon}=\epsilon_{\theta}(\tau^{k},\emptyset,k)+\omega\Big(\epsilon_{\theta}(\tau^{k},y(\tau),k)-\epsilon_{\theta}(\tau^{k},\emptyset,k)\Big),$
where $\emptyset$ denotes a fixed null input token for the condition and $\omega>0$ controls the guidance strength. This yields trajectory samples aligned with task objectives while retaining coverage of the dataset. To train both pathways in a single model, one uses conditional dropout: the condition $y(\tau)$ is randomly masked with probability $p$, controlled by a Bernoulli variable $\beta$:
$\mathcal{L}(\theta)=\mathbb{E}_{k,\;\tau,\;\epsilon,\;\beta\sim\mathrm{Bern}(p)}\Big[\;\|\epsilon-\epsilon_{\theta}(\tau^{k},\beta y(\tau),k)\|^{2}\;\Big].$
Here $\beta=0$ masks the condition and sets it to null token $\emptyset$. This allows training both a conditional and unconditional predictor simultaneously. At inference, the two are recombined via classifier-free guidance (CFG) as in eq. (1).
Figure 2: Illustrative scenario: We compare sampled state trajectories with open-loop simulations obtained by executing the sampled actions on a simple car model (cf. App. C). Diffuser produces infeasible trajectories, and Decision Diffuser yields plausible states whose actions diverge when executed. In contrast, MPDiffuser generates trajectories that are faithful to system dynamics.
@@ 3 Method
Below, we introduce the components of our compositional framework, describe the sampling procedure, and show how it can be used for the constrained decision-making problem.
@@ 3.1 Framework Components
$\mathcal{L}^{\mathrm{pl}}(\theta)=\mathbb{E}_{\tau,k,\epsilon,\beta}\left[\;\big\|[\epsilon_{x},\epsilon_{u}]-\epsilon_{\theta}^{\mathrm{pl}}\!\big(\tau^{k}_{x},\tau_{u}^{k}\mid k,x_{0},\beta y(\tau)\big)\big\|^{2}\;\right].$
@@
Complementing the planner, the dynamics model acts as the feasibility filter—refining state trajectories so that imagined rollouts remain faithful to the system’s underlying physics. In contrast to the planner, which models full state–action rollouts, the dynamics model is a conditional diffusion model over states only, given the initial state $x_{0}$, an action sequence is corrupted by the same noise level as states $\tau_{u}^{k}$, and conditioning vector $y(\tau)$. By treating actions purely as inputs, the model dedicates its capacity to capturing state transitions, thereby yielding sharper dynamics consistency than the planner. We train our dynamics diffusion model $\epsilon_{\vartheta}^{\mathrm{dyn}}(\tau_{x}^{k}\mid\tau_{u}^{k},k,x_{0},y(\tau))$ using:
$\mathcal{L}^{\mathrm{dyn}}(\vartheta)=\mathbb{E}_{\tau,k,\epsilon,\beta}\left[\;\big\|\epsilon_{x}-\epsilon_{\vartheta}^{\mathrm{dyn}}\!\big(\tau_{x}^{k}\mid\tau_{u}^{k},k,x_{0},\beta y(\tau)\big)\big\|^{2}\;\right].$
@@
The ranker is a practical task-aware module that evaluates sampled trajectories against desired criteria, selecting rollouts that achieve high reward, respect safety budgets. Within our framework, it is treated as a flexible scoring function $\rho(\tau)$ that assigns preference values to trajectories. This allows incorporation of both domain knowledge and data-driven objectives. Formally, given sampled trajectories $\{\tau_{j}\}$, the ranker outputs $\tau^{\star}=\arg\max_{\tau_{j}}\rho(\tau_{j})$, with $\rho$ defined by the task e.g. maximizing return under constraints, or minimizing goal distance. This design balances flexibility and structure: when objectives are clearly specified, $\rho$ can be explicitly defined analytically, while in settings with implicit preferences, $\rho$ may be learned from data.